# On Edge Coloring of Multigraphs

### v4

@@ 5.2.2 Case Verification
Since $V(T)$ is not elementary, there exist a color $\alpha\in\overline{\varphi}(y_{p})\cap\overline{\varphi}(v)$ for some $v\in V(T(y_{p-1}))$. Note that $\gamma_{n}\notin\Gamma^{q}\cup D_{n,q}$; we use this fact many times in our proof without mentioning it for a clean presentation. We assume without loss of generality, one of $\gamma$ and $\tau$ in $\overline{\varphi}(x_{e})$, say $\gamma$, is different from $\gamma_{n}$.
@@
Assume $\alpha=\gamma_{n}$. Then we just let $\pi=\varphi/P_{y_{p}}(\gamma_{n},\gamma,\varphi)$. By Claim 5.7 (i), $T_{n,q}$ is an ETT with a hierarchy of $(q-1)$-levels up to itself under the $T_{n}+f_{n}$ mod $\varphi$ coloring $\pi$. Moreover, since $\gamma_{n},\gamma\in\overline{\varphi}(T_{n})$ and $\gamma_{n},\gamma\notin\Gamma^{q}$ following Definition 4, it is routine to see that $T$ remains an ETT under $\pi$ with a good hierarchy of $q$-levels and $\pi\in\overline{\varphi}(y_{p})\cap\overline{\varphi}(T_{n})$, and therefore $T$ remains a minimum counterexample under $\pi$. So the claim is established by replacing $\varphi$ with $\pi$. ∎
@@
In this case, we have $\alpha\in\overline{\varphi}(T_{n})$ and $\alpha$ is closed for $T_{n}$. Since $\gamma_{n}$ is also closed for $T_{n}$ under $\varphi$, $P_{y_{1}}(\alpha,\gamma_{n},\varphi)$ is different from $P_{v(\alpha)}(\alpha,\gamma_{n},\varphi)=P_{v(\beta)}(\alpha,\gamma_{n},\varphi)$ by Claim 5.6. Let $\pi=\varphi/P_{y_{1}}(\alpha,\gamma_{n},\varphi)$. By Claim 5.7 (i), $\pi$ is $T_{n}+f_{n}$ mod $\varphi$, and therefore $T_{n}+f_{n}$ is an ETT with $n$-rungs under $\pi$ with the same extension types, connecting colors, companion colors and connecting edges as under $\varphi$. Let $h$ be the initial index for $\Theta_{n}$. Consequently, $y_{1}$ belongs to a $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ following Algorithm 3.1 ($h=n$ when $\Theta_{n}$=IE). On the other hand, since $\gamma_{n}=\gamma_{h}\in\overline{\pi}(y_{1})$, $f_{n}$ is no longer an edge contained in a $(\gamma_{h},\delta_{h})$-ear of $T_{h}$, a contradiction.
@@
In this case $e_{1}$ is colored by $\beta\in\Gamma^{q-1}$. So $\beta\notin\{\gamma_{n},\gamma\}$ and $\beta\in\overline{\varphi}(T_{n,q})$ by Definition 4. Note that $\alpha\neq\gamma_{n}$ following Claim 5.8. Since $\gamma$ is closed for $T_{n,q}$ by Claim 5.1 and $\gamma\neq\gamma_{n}$, $P_{x_{e}}(\gamma,\alpha,\varphi)=P_{v(\alpha)}(\gamma,\alpha,\varphi)$ is the only $(\gamma,\alpha)$-path intersecting $T_{n,q}$ by Claim 5.6. Thus $P_{y_{1}}(\gamma,\alpha,\varphi)$ is different from $P_{x_{e}}(\gamma,\alpha,\varphi)$. Following Claim 5.7, $T_{n,q}$ is an ETT with a good hierarchy up to itself of $(q-1)$-levels under $\pi=\varphi/P_{y_{1}}(\gamma,\alpha,\varphi)$. Moreover, because $\beta\notin\{\gamma_{n},\gamma\}$, $\pi$ is $(T,D_{n},\varphi)$-weakly stable, and therefore it remains a counterexample under $\pi$. However, since $y_{1}$, $x_{e}$, and $v(\beta)$ induce at least two $(\beta,\gamma)$-path intersecting $T_{n,q}$ under $\pi$ ($v(\beta)=x_{e}$ may happen), we reach a contradiction to Claim 5.6 under $\pi$. ∎
Recall that the path number $p(T)$ of $T$ is the smallest index $i\in{1,...,p}$ such that the sequence $y_{i}T:=(y_{i},e_{i+1},...,e_{p},y_{p})$ corresponds to path in $G$, where $p\geq 2$ by Claim 5.9. Depending on the value of $p(T)$, we divide the remainder of the proof into three cases. In these three cases, we are going to perform a lot of Kempe changes based on Claims 5.6 and 5.7 similarly as above, where Claim 5.6 is used to show that a path $P$ is different from $P_{v(\alpha)}(\alpha,\beta,\varphi)$ to set up Claim 5.7. We may skip mentioning some of these details when they are straightforward. By Claim 5.8, when applying Claims 5.6 and 5.7 to $\alpha$ later, we will assume $\alpha\neq\gamma_{n}$ by default without mentioning Claim 5.8, unless specified otherwise.
@@
We may assume $\alpha\in\overline{\varphi}(y_{i})\cap\overline{\varphi}(y_{p})$ for some $11\leq i\leq p-$ 1.
@@
Now we assume that $j=1$. In this case $q>0$, since when $q=0$ we have $\alpha=\delta_{n}\notin\overline{\varphi}(T_{n})$. Therefore, $\alpha=\varphi(e_{1})$ where $\alpha\in\Gamma^{q-1}$ following Definition 4 (iv). Note that then $\alpha\notin\Gamma^{q}$ by Definition 4 (i). Note that we also have $\gamma\notin\Gamma^{q}$ by Definition 4 (i). Let $\pi=\varphi/P_{y_{p}}(\alpha,\gamma,\varphi)$ Following Claim 5.6 and Claim 5.7 (i), since $\alpha,\gamma\notin\Gamma^{q}$ and $\alpha,\gamma\in\overline{\varphi}(T_{n,q})$ with $\gamma$ closed for $T_{n,q}$, we have $P_{x_{e}}(\alpha,\gamma,\varphi)=P_{v(\alpha)}(\alpha,\gamma,\varphi)$, and it is routine to check that $T$ remains a counterexample with a good hierarchy under $\pi$ and $\pi(e_{1})=\varphi(e_{1})=\alpha$. But now either $\gamma\notin\pi(T-T_{n,q})$, or $\gamma\in\pi(T-T_{n,q})$ but the first edge colored by $\gamma$ is not $e_{1}$, we reach previous cases. ∎
@@
Following Claim 5.10, we assume that $i<p-1$ but $i$ is already maximum among all counterexamples. Let $\beta\in\overline{\varphi}(y_{i+1})$. So $\beta\neq\gamma_{n}$. We first consider the case $\alpha,\beta\notin D_{n,q}$. Then $\alpha,\beta\notin\overline{\varphi}(T_{n,q})\cup D_{n,q}$. Since $y_{i-1}$ is an end of $e_{i}$, we have $\varphi(e_{i})\neq\alpha$. So $\alpha\notin\varphi(T(y_{i})-T_{n,q})$ by TAA. Thus by Claim 5.6, $P_{y_{p}}(\alpha,\beta,\varphi)$ is different from $P_{y_{i}}(\alpha,\beta,\varphi)=P_{y_{i+1}}(\alpha,\beta,\varphi)$. Therefore following Claim 5.7, $T$ is a counterexample with a good hierarchy under $\pi=\varphi/P_{y_{p}}(\alpha,\beta,\varphi)$. But $\beta\in\overline{\pi}(y_{i+1})\cap\overline{\pi}(y_{p})$, a contradiction to the maximality of $i$.
@@
Note that by (1), following Claim 5.2 and Remark 3, we have a smaller counterexample $T(y_{p-1})$ with a good hierarchy of $q$-levels, contradicting our assumption on $T$.
@@
(2) Under $\varphi$, all the $\delta_{h}$-colored boundary edges of $T_{h}$ belong to some $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ except for the only $(\gamma_{h},\delta_{h})$-exit edge.
Since $\pi$ is also $(T_{h},D_{h},\varphi)$-stable and $T_{h}$ is $RE$ finished, again by (A2),
[…]
@@
$\bullet$ $T^{*}\hskip 1.70717pt=(T_{n,q}^{*},e_{1},y_{1},e_{2},\ldots,y_{p-2},e_{p},y_{p},e_{p-1},y_{p-1})$. Note that $T^{-}$ is obtained from $T$ by deleting $y_{p-1}$ and $T^{*}$ arises from $T$ by interchanging the order of $(e_{p-1},y_{p-1})$ and $(e_{p},y_{p})$.
@@
We first consider the case $\theta\notin\Gamma^{q}$. In this case, it is easy to check that $T^{*}$ is also a counterexample with a good hierarchy of $q$-levels under $\varphi$. Following (1), we have $\gamma_{m1},\gamma_{m2},\delta_{m}\notin\varphi(T^{*}-T_{n,q}-f_{n})$. Thus by Lemma 5.6 on $T^{*}$, we have $P_{v(\gamma_{mj})}(\delta_{m},\gamma_{mj},\varphi)=P_{y_{p}}(\delta_{m},\gamma_{mj},\varphi)$ for $j=1,2$, a contradiction to (2).
Now we assume $\theta\in\Gamma^{q}$. Without loss of generality, we say $\theta=\gamma_{k1}$ with $\delta_{k}\in D_{n,q}$. Note that if $\delta_{k}\notin\overline{\varphi}(y_{p-1})$, then $T^{*}$ is also a counterexample with a good hierarchy of $q$-levels under $\varphi$, and therefore we can get a contradiction to (2) again as above. So we may assume $\delta_{k}\in\overline{\varphi}(y_{p-1})$. So following Definition 4 (i),
@@
Note that in this case, $T^{*}$ is also a counterexample with a good hierarchy, unless $\theta\in\Gamma^{q}_{m}$ for some $\delta_{m}\in D_{n,q}$ with $\delta_{m}\in\overline{\varphi}(y_{p-1})$. We first assume that there does not exist such $m$ that $\theta\in\Gamma^{q}_{m}$ with $\delta_{m}\in\ D_{n,q}$ and $\delta_{m}\in\overline{\varphi}(y_{p-1})$. By Claim 5.5, we have $|\overline{\varphi}(T(y_{p-2}))\setminus\{\gamma_{n}\}-\varphi(T(y_{p-2})-T_{n,q})|\geq 2n+2$. Recall that $|D_{n,q}|\leq n$ and $n\geq 1$. Since are only two edges in $T-T(y_{p-2})$, there exists a color $\beta\in\overline{\varphi}(T(y_{p-2}))-D_{n,q}\cup\{\gamma_{n}\}$ such that $\beta\notin\varphi(T-T_{n,q})$. Note that $\beta\notin\overline{\varphi}(y_{p})$. Otherwise, $T^{-}$ is a smaller counterexample, giving a contradiction to the minimum assumption of $T$. Now since $\alpha,\beta\notin D_{n,q}$, $\alpha,\beta\neq\gamma_{n}$ and $\beta\notin\varphi(T-T_{n,q})$, we have $P_{v(\beta)}(\alpha,\beta,\varphi)=P_{y_{p-1}}(\alpha,\beta,\varphi)$ by applying Claim 5.6 on $T$. On the other hand, by applying Claim 5.6 on $T^{*}$, we see that $P_{v(\beta)}(\alpha,\beta,\varphi)=P_{y_{p}}(\alpha,\beta,\varphi)$, a contradiction.
@@
    Under $\varphi$, there exists a non-elementary ETT $T^{\prime}$ with a good hierarchy of $q$-levels but $p(T^{\prime})<p(T)$.
@@
Clearly $T^{\prime}$ is an ETT under $\varphi$ as it can by obtained by TAA from $T_{n,q}$, and $T_{n,q}$ has a good hierarchy of $(q-1)$-levels up to itself. Note that $p(T^{\prime})<P(T)$. We proceed with the proof by considering three cases: $\alpha\notin\Gamma^{q}$, $\alpha\in\Gamma^{q}-\varphi(T-T_{n,q})$ and $\alpha\in\Gamma^{q}\cap\varphi(T-T_{n,q})$.
Case I: $\alpha\notin\Gamma^{q}$.
Since $\alpha,\beta\in\overline{\varphi}(T_{y_{p-2}})$ and if $\beta\in\Gamma^{q}_{m}$ then $\delta_{m}\in\overline{\varphi}(T(y_{p-2}))$, we see that $T^{\prime}=T_{n,q+1}$ satisfies Definition 4 (i). Hence it is routine to see that statement (b) holds, a contradiction.
Case II: $\alpha\in\Gamma^{q}\cap\varphi(T-T_{n,q})$.
Assume $\alpha=\gamma_{k1}\in\Gamma^{q}$ with $\delta_{k}\in D_{n,q}$. Since $\varphi(e_{p})\neq\alpha$, $\alpha\in\varphi(T_{y_{p-1}}-T_{n,q})$. Therefore we must have $\delta_{m}\in\overline{\varphi}(T(y_{p-2}))$ following Definition 4 (i). Since $\delta_{m}\in\overline{\varphi}(T(y_{p-2}))$ when $\beta\in\Gamma^{q}_{m}$, we again see that $T^{\prime}=T_{n,q+1}$ satisfies Definition 4 (i). Hence it is routine to see that statement (b) holds, a contradiction.
Case III: $\alpha\in\Gamma^{q}-\varphi(T-T_{n,q})$.
Then $\alpha,\beta\notin\varphi(T-T_{n,q})$ and $\alpha,\beta\neq\gamma_{n}$. Let $\pi=\varphi/P$. Again following Claims 5.6 and 5.7, it is routine to check that $T$ remains a counterexample with a good hierarchy, and statement (a) holds. ∎
@@
$12\leq p(T)\leq p-$ 1.
Define $I_{\varphi}=\{i:\overline{\varphi}(y_{p})\cap\overline{\varphi}(y_{i})\neq\emptyset\}$. So $I_{\varphi}=\emptyset$ when $\{v:\overline{\varphi}(y_{p})\cap\overline{\varphi}(v)\neq\emptyset\}\subseteq V(T_{n,q})$. For convention, we denote $\max{I_{\varphi}}=0$ when $I_{\varphi}=\emptyset$. Let $j=p(T)$. By the assumption of the current case, we have $j\geq 2$ and $e_{j}\notin E(y_{j-1},y_{j})$. We first consider the case $max(I_{\varphi})\geq j$. This case is similar to Case 1 and can be handled in the same fashion: We first show that $\max(I_{\varphi})=p-1$ the same was as in Claim 5.11, and replace color $\varphi(e_{p})$ by $\alpha$ to get a smaller counterexample with a new color $\pi$. The proof of $\pi$ being $T_{n}+f_{n}$ is also the same as the end in Case 1. We omit the details.
@@
Following Claim 5.8, let $\alpha\in\overline{\varphi}(y_{p})\cap\overline{\varphi}(T(y_{j-1}))$ with $\alpha\neq\gamma_{n}$. We consider the following two cases.
Case I
$\alpha\in\overline{\varphi}(y_{j-1})$.
@@
Case II
$\alpha\in\overline{\varphi}(T(y_{j-2}))$.
@@
Now we assume that $\delta_{m}\in\overline{\varphi}(T(y_{p-1}))$. Note that $\delta_{m}\notin\overline{\varphi}(T_{n,q})$ since $\delta_{m}\in D_{n,q}$. So $\delta_{m}\in\overline{\varphi}(y_{k})$ for some $11\leq k\leq p-$ 1. If $k<j-1$, then we have Claim 5.13 already, so we assume $k\geq j-1$. Following Definition 4 (i), we have $\gamma_{m1}\notin\varphi(T(y_{k})-T_{n,q})$. So following Claims 5.6 and 5.7, it is routine to check that $T$ is a counterexample with a good hierarchy under $\pi=\varphi/P_{y_{p}}(\delta_{m},\gamma_{m1},\varphi)$ with $\delta_{m}\in\overline{\pi}(y_{p})\cap\overline{\pi}(y_{k})$. Now if $k=j-1$, we return to Case I. Otherwise, $k\geq j$ and we reach the case that $max(I_{\varphi})\geq j$, and then proceed as in Case 1. ∎
@@
Clearly $T^{\prime}$ is an ETT under $\varphi$ as it can by obtained by TAA from $T_{n,q}$, and $T_{n,q}$ has a good hierarchy of $(q-1)$-levels up to itself. By the assumptions on $\alpha$ and $\beta$, $T^{\prime}$ also satisfies Definition 4 (i). So following Remark 3, $T^{\prime}$ is a counterexample under $\varphi$. Note that we have $p(T^{\prime})<P(T)$, giving a contradiction to the minimum assumption on $T$.

### v5

@@ 5.2.2 Case Verification
Since $V(T)$ is not elementary, there exist a color $\alpha\in\overline{\varphi}(y_{p})\cap\overline{\varphi}(v)$ for some $v\in V(T(y_{p-1}))$. Note that $\gamma_{n}\notin\Gamma^{q}\cup D_{n,q}$; we use this fact many times in our proof without mentioning it for a clean presentation. Recall that for simplification of notations, we let $\Gamma^{q}_{m}=\{\gamma_{m1},\gamma_{m2}\}$ for $\delta_{m}\in D_{n,q}$, and reserve $\gamma,\tau$ for the two colors in $\overline{\varphi}(x_{e})$ in Definition 4. We assume without loss of generality, one of $\gamma$ and $\tau$ in $\overline{\varphi}(x_{e})$, say $\gamma$, is different from $\gamma_{n}$.
@@
Assume $\alpha=\gamma_{n}$. Then we just let $\pi=\varphi/P_{y_{p}}(\gamma_{n},\gamma,\varphi)$. By Claim 5.7 (i), $T_{n,q}$ is an ETT with a hierarchy of (q1)(q-1) levels up to itself under the $T_{n}+f_{n}$ mod $\varphi$ coloring $\pi$. Moreover, since $\gamma_{n},\gamma\in\overline{\varphi}(T_{n})$ and $\gamma_{n},\gamma\notin\Gamma^{q}$ following Definition 4, it is routine to see that $T$ remains an ETT under $\pi$ with a good hierarchy of $q$ levels and $\pi\in\overline{\varphi}(y_{p})\cap\overline{\varphi}(T_{n})$, and therefore $T$ remains a minimum counterexample under $\pi$. So the claim is established by replacing $\varphi$ with $\pi$. ∎
@@
In this case, we have $\alpha\in\overline{\varphi}(T_{n})$ and $\alpha$ is closed for $T_{n}$. Since $\gamma_{n}$ is also closed for $T_{n}$ under $\varphi$, $P_{y_{1}}(\alpha,\gamma_{n},\varphi)$ is different from $P_{v(\alpha)}(\alpha,\gamma_{n},\varphi)=P_{v(\beta)}(\alpha,\gamma_{n},\varphi)$ by Claim 5.6. Let $\pi=\varphi/P_{y_{1}}(\alpha,\gamma_{n},\varphi)$. By Claim 5.7 (i), $\pi$ is $T_{n}+f_{n}$ mod $\varphi$, and therefore $T_{n}+f_{n}$ is an ETT with $n$-rungs under $\pi$ with the same extension types, connecting colors, companion colors and connecting edges as under $\varphi$. Let $h$ be the initial index for $\Theta_{n}$. Consequently, $y_{1}$ belongs to a $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ following Algorithm 3.1 ($h=n$ when $\Theta_{n}=\operatorname{IE}$). On the other hand, since $\gamma_{n}=\gamma_{h}\in\overline{\pi}(y_{1})$, $f_{n}$ is no longer an edge contained in a $(\gamma_{h},\delta_{h})$-ear of $T_{h}$, a contradiction.
@@
In this case $e_{1}$ is colored by $\beta\in\Gamma^{q-1}$. So $\beta\notin\{\gamma_{n},\gamma\}$ and $\beta\in\overline{\varphi}(T_{n,q})$ by Definition 4. Note that $\alpha\neq\gamma_{n}$ following Claim 5.8. Since $\gamma$ is closed for $T_{n,q}$ by Claim 5.1 and $\gamma\neq\gamma_{n}$, $P_{x_{e}}(\gamma,\alpha,\varphi)=P_{v(\alpha)}(\gamma,\alpha,\varphi)$ is the only $(\gamma,\alpha)$-path intersecting $T_{n,q}$ by Claim 5.6. Thus $P_{y_{1}}(\gamma,\alpha,\varphi)$ is different from $P_{x_{e}}(\gamma,\alpha,\varphi)$. Following Claim 5.7, $T_{n,q}$ is an ETT with a good hierarchy up to itself of (q1)(q-1) levels under $\pi=\varphi/P_{y_{1}}(\gamma,\alpha,\varphi)$. Moreover, because $\beta\notin\{\gamma_{n},\gamma\}$, $\pi$ is $(T,D_{n},\varphi)$-weakly stable, and therefore it remains a counterexample under $\pi$. However, since $y_{1}$, $x_{e}$, and $v(\beta)$ induce at least two $(\beta,\gamma)$-paths intersecting $T_{n,q}$ under $\pi$ ($v(\beta)=x_{e}$ may happen), we reach a contradiction to Claim 5.6 under $\pi$. ∎
Recall that the path number $p(T)$ of $T$ is the smallest index $i\in\{1,...,p\}$ such that the sequence $y_{i}T:=(y_{i},e_{i+1},...,e_{p},y_{p})$ corresponds to a path in $G$, where $p\geq 2$ by Claim 5.9. Depending on the value of $p(T)$, we divide the remainder of the proof into three cases. In these three cases, we are going to perform a lot of Kempe changes based on Claims 5.6 and 5.7 similarly as above, where Claim 5.6 is used to show that a path $P$ is different from $P_{v(\alpha)}(\alpha,\beta,\varphi)$ to set up Claim 5.7. We may skip mentioning some of these details when they are straightforward. By Claim 5.8, when applying Claims 5.6 and 5.7 to $\alpha$ later, we will assume $\alpha\neq\gamma_{n}$ by default without mentioning Claim 5.8, unless specified otherwise.
@@
We may assume $\alpha\in\overline{\varphi}(y_{i})\cap\overline{\varphi}(y_{p})$ for some $1\leq i\leq p-1$.
@@
Now we assume that $j=1$. In this case $q>0$, since when $q=0$ we have $\alpha=\delta_{n}\notin\overline{\varphi}(T_{n})$. Therefore, $\alpha=\varphi(e_{1})$ where $\alpha\in\Gamma^{q-1}$ following Definition 4 (iv). Note that then $\alpha\notin\Gamma^{q}$ by Definition 4 (i). Note that we also have $\gamma\notin\Gamma^{q}$ by Definition 4 (i). Let $\pi=\varphi/P_{y_{p}}(\alpha,\gamma,\varphi)$. Following Claim 5.6 and Claim 5.7 (i), since $\alpha,\gamma\notin\Gamma^{q}$ and $\alpha,\gamma\in\overline{\varphi}(T_{n,q})$ with $\gamma$ closed for $T_{n,q}$, we have $P_{x_{e}}(\alpha,\gamma,\varphi)=P_{v(\alpha)}(\alpha,\gamma,\varphi)$, and it is routine to check that $T$ remains a counterexample with a good hierarchy under $\pi$ and $\pi(e_{1})=\varphi(e_{1})=\alpha$. But now either $\gamma\notin\pi(T-T_{n,q})$, or $\gamma\in\pi(T-T_{n,q})$ but the first edge colored by $\gamma$ is not $e_{1}$, we reach previous cases. ∎
@@
Following Claim 5.10, we assume that i<p1i<p-1 but $i$ is already maximum among all counterexamples. Let $\beta\in\overline{\varphi}(y_{i+1})$. So $\beta\neq\gamma_{n}$. We first consider the case $\alpha,\beta\notin D_{n,q}$. Then $\alpha,\beta\notin\overline{\varphi}(T_{n,q})\cup D_{n,q}$. Since $y_{i-1}$ is an end of $e_{i}$, we have $\varphi(e_{i})\neq\alpha$. So $\alpha\notin\varphi(T(y_{i})-T_{n,q})$ by TAA. Thus by Claim 5.6, $P_{y_{p}}(\alpha,\beta,\varphi)$ is different from $P_{y_{i}}(\alpha,\beta,\varphi)=P_{y_{i+1}}(\alpha,\beta,\varphi)$. Therefore following Claim 5.7, $T$ is a counterexample with a good hierarchy under $\pi=\varphi/P_{y_{p}}(\alpha,\beta,\varphi)$. But $\beta\in\overline{\pi}(y_{i+1})\cap\overline{\pi}(y_{p})$, a contradiction to the maximality of $i$.
@@
Note that by (1), following Claim 5.2 and Remark 4, we have a smaller counterexample $T(y_{p-1})$ with a good hierarchy of $q$ levels, contradicting our assumption on $T$.
@@
(2) Under $\varphi$, all the $\delta_{h}$ colored boundary edges of $T_{h}$ belong to some $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ except for the only $(\gamma_{h},\delta_{h})$-exit edge.
Since $\pi$ is also $(T_{h},D_{h},\varphi)$-stable and $T_{h}$ is $\operatorname{RE}$ finished, again by (A2),
(3) $T_{h}$ is an ETT with (h1)(h-1)-rungs under $\pi$.
@@
$\bullet$ $T^{*}\thinspace=(T_{n,q}^{*},e_{1},y_{1},e_{2},\ldots,y_{p-2},e_{p},y_{p},e_{p-1},y_{p-1})$. Note that $T^{-}$ is obtained from $T$ by deleting $y_{p-1}$ and $T^{*}$ arises from $T$ by interchanging the order of $(e_{p-1},y_{p-1})$ and $(e_{p},y_{p})$.
@@
We first consider the case $\theta\notin\Gamma^{q}$. In this case, it is easy to check that $T^{*}$ is also a counterexample with a good hierarchy of $q$ levels under $\varphi$. Following (1), we have $\gamma_{m1},\gamma_{m2},\delta_{m}\notin\varphi(T^{*}-T_{n,q}-f_{n})$. Thus by Lemma 5.6 on $T^{*}$, we have $P_{v(\gamma_{mj})}(\delta_{m},\gamma_{mj},\varphi)=P_{y_{p}}(\delta_{m},\gamma_{mj},\varphi)$ for $j=1,2$, a contradiction to (2).
Now we assume $\theta\in\Gamma^{q}$. Without loss of generality, we say $\theta=\gamma_{k1}$ with $\delta_{k}\in D_{n,q}$. Note that if $\delta_{k}\notin\overline{\varphi}(y_{p-1})$, then $T^{*}$ is also a counterexample with a good hierarchy of $q$ levels under $\varphi$, and therefore we can get a contradiction to (2) again as above. So we may assume $\delta_{k}\in\overline{\varphi}(y_{p-1})$. So following Definition 4 (i),
@@
Note that in this case, $T^{*}$ is also a counterexample with a good hierarchy, unless $\theta\in\Gamma^{q}_{m}$ for some $\delta_{m}\in D_{n,q}$ with $\delta_{m}\in\overline{\varphi}(y_{p-1})$. We first assume that there does not exist such $m$ that $\theta\in\Gamma^{q}_{m}$ with $\delta_{m}\in\ D_{n,q}$ and $\delta_{m}\in\overline{\varphi}(y_{p-1})$. By Claim 5.5, we have $|(\overline{\varphi}(T(y_{p-2}))\setminus\{\gamma_{n}\})\setminus\varphi(T(y_{p-2})-T_{n,q})|\geq 2n+2$. Recall that $|D_{n,q}|\leq n$ and $n\geq 1$. Since there are only two edges in $T-T(y_{p-2})$, there exists a color $\beta\in\overline{\varphi}(T(y_{p-2}))\setminus(D_{n,q}\cup\{\gamma_{n}\})$ such that $\beta\notin\varphi(T-T_{n,q})$. Note that $\beta\notin\overline{\varphi}(y_{p})$. Otherwise, $T^{-}$ is a smaller counterexample, giving a contradiction to the minimum assumption of $T$. Now since $\alpha,\beta\notin D_{n,q}$, $\alpha,\beta\neq\gamma_{n}$ and $\beta\notin\varphi(T-T_{n,q})$, we have $P_{v(\beta)}(\alpha,\beta,\varphi)=P_{y_{p-1}}(\alpha,\beta,\varphi)$ by applying Claim 5.6 on $T$. On the other hand, by applying Claim 5.6 on $T^{*}$, we see that $P_{v(\beta)}(\alpha,\beta,\varphi)=P_{y_{p}}(\alpha,\beta,\varphi)$, a contradiction.
@@
    Under $\varphi$, there exists a non-elementary ETT $T^{\prime}$ with a good hierarchy of $q$ levels but $p(T^{\prime})<p(T)$.
@@
Clearly $T^{\prime}$ is an ETT under $\varphi$ as it can be obtained by TAA from $T_{n,q}$, and $T_{n,q}$ has a good hierarchy of (q1)(q-1) levels up to itself. Note that $p(T^{\prime})<p(T)$. We proceed with the proof by considering three cases: $\alpha\notin\Gamma^{q}$, $\alpha\in\Gamma^{q}-\varphi(T-T_{n,q})$ and $\alpha\in\Gamma^{q}\cap\varphi(T-T_{n,q})$. Case I: $\alpha\notin\Gamma^{q}$. Since $\alpha,\beta\in\overline{\varphi}(T_{y_{p-2}})$ and if $\beta\in\Gamma^{q}_{m}$ then $\delta_{m}\in\overline{\varphi}(T(y_{p-2}))$, we see that $T^{\prime}=T_{n,q+1}$ satisfies Definition 4 (i). Hence it is routine to see that statement (b) holds, a contradiction.
Case II: $\alpha\in\Gamma^{q}\cap\varphi(T-T_{n,q})$. Assume $\alpha=\gamma_{k1}\in\Gamma^{q}$ with $\delta_{k}\in D_{n,q}$. Since $\varphi(e_{p})\neq\alpha$, $\alpha\in\varphi(T_{y_{p-1}}-T_{n,q})$. Therefore we must have $\delta_{m}\in\overline{\varphi}(T(y_{p-2}))$ following Definition 4 (i). Since $\delta_{m}\in\overline{\varphi}(T(y_{p-2}))$ when $\beta\in\Gamma^{q}_{m}$, we again see that $T^{\prime}=T_{n,q+1}$ satisfies Definition 4 (i). Hence it is routine to see that statement (b) holds, a contradiction.
Case III: $\alpha\in\Gamma^{q}-\varphi(T-T_{n,q})$. Then $\alpha,\beta\notin\varphi(T-T_{n,q})$ and $\alpha,\beta\neq\gamma_{n}$. Let $\pi=\varphi/P$. Again following Claims 5.6 and 5.7, it is routine to check that $T$ remains a counterexample with a good hierarchy, and statement (a) holds. ∎
@@
$2\leq p(T)\leq p-1$.
Define $I_{\varphi}=\{i:\overline{\varphi}(y_{p})\cap\overline{\varphi}(y_{i})\neq\emptyset\}$. So $I_{\varphi}=\emptyset$ when $\{v:\overline{\varphi}(y_{p})\cap\overline{\varphi}(v)\neq\emptyset\}\subseteq V(T_{n,q})$. For convention, we denote $\max{I_{\varphi}}=0$ when $I_{\varphi}=\emptyset$. Let $j=p(T)$. By the assumption of the current case, we have $j\geq 2$ and $e_{j}\notin E(y_{j-1},y_{j})$. We first consider the case $max(I_{\varphi})\geq j$. This case is similar to Case 1 and can be handled in the same fashion: We first show that $\max(I_{\varphi})=p-1$ the same way as in Claim 5.11, and replace color $\varphi(e_{p})$ by $\alpha$ to get a smaller counterexample with a new color $\pi$. The proof of $\pi$ being $T_{n}+f_{n}$ is also the same as the end in Case 1. We omit the details.
@@
Following Claim 5.8, let $\alpha\in\overline{\varphi}(y_{p})\cap\overline{\varphi}(T(y_{j-1}))$ with $\alpha\neq\gamma_{n}$. We consider the following two cases. Case I $\alpha\in\overline{\varphi}(y_{j-1})$.
@@
Case II $\alpha\in\overline{\varphi}(T(y_{j-2}))$.
@@
Now we assume that $\delta_{m}\in\overline{\varphi}(T(y_{p-1}))$. Note that $\delta_{m}\notin\overline{\varphi}(T_{n,q})$ since $\delta_{m}\in D_{n,q}$. So $\delta_{m}\in\overline{\varphi}(y_{k})$ for some $1\leq k\leq p-1$. If k<j1k<j-1, then we have Claim 5.13 already, so we assume $k\geq j-1$. Following Definition 4 (i), we have $\gamma_{m1}\notin\varphi(T(y_{k})-T_{n,q})$. So following Claims 5.6 and 5.7, it is routine to check that $T$ is a counterexample with a good hierarchy under $\pi=\varphi/P_{y_{p}}(\delta_{m},\gamma_{m1},\varphi)$ with $\delta_{m}\in\overline{\pi}(y_{p})\cap\overline{\pi}(y_{k})$. Now if k=j1k=j-1, we return to Case I. Otherwise, $k\geq j$ and we reach the case that $max(I_{\varphi})\geq j$, and then proceed as in Case 1. ∎
@@
Clearly $T^{\prime}$ is an ETT under $\varphi$ as it can by obtained by TAA from $T_{n,q}$, and $T_{n,q}$ has a good hierarchy of (q1)(q-1) levels up to itself. By the assumptions on $\alpha$ and $\beta$, $T^{\prime}$ also satisfies Definition 4 (i). So following Remark 4, $T^{\prime}$ is a counterexample under $\varphi$. Note that we have $p(T^{\prime})<p(T)$, giving a contradiction to the minimum assumption on $T$.